OpenStudy (anonymous):

Find the solution for the equation 7x + 35 = -3x + 5

pooja195 (pooja195):

Start by adding 3x to both sides 7+3=?

OpenStudy (1andrewalarcon1):

You have to isolate x first on the left side just like pooja said

OpenStudy (anonymous):

7x + 35 = -3x + 5 +3x - 5 = +3x - 5 ----------------------- 10x +30

pooja195 (pooja195):

No.. 10x+35=5 Subtract 35 from both sides we want to isolate x

OpenStudy (anonymous):

so what is the answer?

OpenStudy (1andrewalarcon1):

Try subtracting 35 instead of 5

pooja195 (pooja195):

You tell me. 5-35=?

OpenStudy (anonymous):

10x-30?

pooja195 (pooja195):

10x=-30 We are isolating x we want to separate the 2 not put them together Now divide both sides by 10 -30/10=?

OpenStudy (anonymous):

x=-3

OpenStudy (1andrewalarcon1):

correct :)

OpenStudy (1andrewalarcon1):

good job

pooja195 (pooja195):

Well done.